# Parameterization-driven Neural Implicit Surfaces Editing

### v2

@@ Appendix C Geometry Editing via Modifying the Parametric Domain $\mathcal{D}$
In addition to the 2D displacement map, we also provide a geometry-editing method (See Figure 16) based on the parametric domain. As shown in Figure 17, we remove the handle of the cup in the parametric domain, and then we render the cup model in volume rendering without any training again.
Figure 17: Editing the geometry of the parametric domain $\mathcal{D}$ naturally induces a modified neural implicit surface. We edit the geometry of the cup by removing and deforming the handle in $\mathcal{D}$. The top row shows the original parametric domain $\mathcal{D}$ and two modified domains. The bottom row shows the neural implicit surface $\mathcal{S}$ corresponded to each parametric domain.
@@
Appendix D Qualitative Comparison
We summarize some parametrization methods (refer to Table 2) and some neural editing methods (refer to Table 3) to illustrate the effectiveness and superiority of our approach. As shown in Table 1, we categorize these texture editing methods with other editing methods into three groups: scene-level, object-level, and pixel-level editing. Scene-level editing methods focus on the entire appearance of a scene including lighting and material. Object-level editing methods employ distinct part-based latent codes to manipulate different attributes of a scene, such as hairstyles in i3DMM [63]. Pixel-level editing, on the other hand, offers a fine-grained editing result by taking into account precise user guidance.
NeuTex [53] Density Init. UV-inv V.R. NA Entangled Sphere
ISO-UV [39] SDF Init. UV D.R. Jacobian Entangled $\mathbb{R}^{2}$
NeP [27] Density Mesh & UV V.R. Angle Disentangled $\mathbb{R}^{2}$
Ours SDF None V.R. Laplace Disentangled Sphere & polycube
Method Proxy Deformation Removal Texture Retraining Category
NeRF-Editing [64, 65] Mesh Y N N N N
NeuMesh [57] Mesh Y N Y Y N
DRFC [55] Cage Y N N N N
Ours Parametric domain Y Y Y N Y
Table 3: Qualitative comparison with other methods for editing the geometry of neural radiance fields. While all these methods exhibit deformation capabilities, they encounter challenges of removing specific parts, such as a cup’s handle, as illustrated in Figure 17. By editing the parametric domain $\mathcal{D}$, we can immediately obtain the deformed neural implicit surfaces through inference on the modified domain $\mathcal{D}^{\prime}$. Our method is trained on a specific category of objects, such as human heads and fruits. This specialized training enables our approach to transfer textures and geometries among objects in the same category.
GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v3

@@ Appendix 0.B Geometry Editing via Modifying the Parametric Domain $\mathcal{D}$
We also provide a geometry-editing method (See Fig. 11) based on the parametric domain. As shown in Fig. 12, we remove the handle of the cup in the parametric domain, and then we render the cup model in volume rendering without any training again. The parametric shape $\mathcal{D}$ is manually created in the Blender [3] and adjusted using techniques like cutting and the Free Form Deformation. We then fit the deformed parametric surface $\mathcal{D}^{\prime}$ and replace the parametric domain SDF with $\mathcal{D}^{\prime}$ , and then infer the deformed source geometry $\mathcal{S}^{\prime}$ and render novel images in volume rendering.
$\mathcal{D}$ $\mathcal{S}$ $\mathcal{D}^{\prime}$ $\mathcal{S}^{\prime}$ $\mathcal{D}^{\prime}$ $\mathcal{S}^{\prime}$
Figure 12: Editing the geometry of the parametric surface $\mathcal{D}$ naturally induces a modified neural implicit surface. We edit the geometry of the cup by removing and deforming the handle in $\mathcal{D}$ using Free Form Deformation [42] in the Blender [3]. The first two columns show the original parametric surface $\mathcal{D}$ and reconstructed surface $\mathcal{S}$. Then we give two modified parametric surfaces $\mathcal{D}^{\prime}$ and the neural implicit surface $\mathcal{S}$ corresponded to each parametric surface.
Appendix 0.C Qualitative Comparison
We summarize some parametrization methods (refer to Tab. 1) to illustrate the effectiveness and superiority of our approach. As shown in Tab. 2, we categorize these texture editing methods with other editing methods into three groups: scene-level, object-level, and pixel-level editing. Scene-level editing methods focus on the entire appearance of a scene including lighting and material. Object-level editing methods employ distinct part-based latent codes to manipulate different attributes of a scene, such as hairstyles in i3DMM [64]. Pixel-level editing, on the other hand, offers a fine-grained editing result by taking into account precise user guidance.
NeuTex [55] Density Init. UV-inv V.R. NA Entangled Sphere
ISO-UV [40] SDF Init. UV D.R. Jacobian Entangled $\mathbb{R}^{2}$
NeP [27] Density Mesh & UV V.R. Angle Disentangled $\mathbb{R}^{2}$
Ours SDF None V.R. Laplace Disentangled Sphere & polycube